# Offline-online approximation of multiscale eigenvalue problems with random defects

### v1

@@ 4.3 OLOD method
We use Theorem 2.2, (14) of Theorem 2.5 and the stability of $\mathscr{C}$ in the energy norm together with the triangle inequality to estimate the last term as
$\!\!\!\!\!\!\!\!\!\!\!\!\norm{\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}-\widehat{m}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})$ (30)
$=m((1-\widehat{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}},(1-\widehat{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}})-m(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})$
|$\leq\absolutevalue{m(\widehat{\mathscr{C}}\hat{u}^{(l+j)}_{H,\mathfrak{k}},\widehat{\mathscr{C}}\hat{u}^{(l+j)}_{H,\mathfrak{k}})}+\absolutevalue{m(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\widehat{\mathscr{C}}\hat{u}^{(l+j)}_{H,\mathfrak{k}})}+\absolutevalue{m(\widehat{\mathscr{C}}\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})}$ |
@@
$\lesssim\Big{(}\norm{\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+2\norm{\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\Big{)}H^{2}$
@@
$\lesssim\Big{(}\norm{(\mathscr{C}-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+2\norm{(\mathscr{C}-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}$
@@
$\hskip 56.9055pt+\norm{(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}+3\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}\Big{)}H^{2}$
$\lesssim\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}\Big{(}g^{2}(\mathfrak{k})+2g(\mathfrak{k})+3+\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}^{2}\mathfrak{k}^{d}+\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}\Big{)}H^{2}.$ (31)
@@
${\widetilde{a}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})-\lambda^{(l)}}={a((1-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})-\lambda^{(l)}}.$
Since $(1-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}\in V^{ms}_{H,\mathfrak{k}}$, by the construction in (24), we obtain that
$\begin{split}a((1-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})&\leq\norm{(1-\mathscr{C}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\leq\norm{(1-\widehat{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}.\end{split}$
Combining with (23), we arrive at
${\widetilde{a}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})-\lambda^{(l)}}$ $\leq\norm{u^{(l+j)}-\hat{u}^{(l+j),ms}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda^{(l)}\norm{u^{(l+j)}-\hat{u}^{(l+j),ms}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}$
$\hskip 56.9055pt+\norm{(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{(\mathscr{C}-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}$
$\lesssim\norm{u^{(l+j)}-\hat{u}^{(l+j),ms}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{(\mathscr{C}-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}.$ (32)
@@
Using (13), and substituting (30) and (32) in (29), we find the desired estimate as
|$\begin{split}\absolutevalue{\frac{\hat{\lambda}^{(l)}_{H,\mathfrak{k}}-\lambda^{(l)}}{\lambda^{(l)}}}&\lesssim\Big{(}\max\limits_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}\norm{\hat{u}^{(l+j)}_{H}}_{\mathcal{A}}^{2}+\Big{(}H^{2}+g(\mathfrak{k})+\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}\Big{)}^{2}+\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}^{2}\mathfrak{k}^{d}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}\\ &\hskip 56.9055pt+g(\mathfrak{k})\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\Big{(}3+g(\mathfrak{k})+\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}\Big{)}H^{2}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}\\ &\lesssim\Big{(}H^{2}+g(\mathfrak{k})+\Big{(}\max\limits_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}+\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}^{2}\mathfrak{k}^{d}\Big{)}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}^{2}.\end{split}$
The result follows, since $\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}^{2}$, similar to (26), can be bounded above by a constant due to $\widehat{a}(\cdot,\cdot)$-coercivity. ∎
@@ 5 Numerical experiments
In this section, we illustrate our theoretical developments with several numerical experiments. We focus on the OLOD method and its alternate variant with the $s$-values from Section 2.5. For details in the calculation of $E_{T}$ (if desired for error control) and discussions of the run-time complexity, we refer to [MV22, Section 5]. Note that aspects on the implementation of the LOD (without offline-online strategy) are explained in [EHMP19]. Our code is based on [HK19] and [Ver21], and freely available at [Kol24].
@@
In our numerical experiments, we will focus on the approximation of the lowest non-trivial eigenvalue, but emphasize that the theory of Section 4 is also valid for multiple and/or higher eigenvalues. By our focus on the lowest eigenvalue, we circumvent any possible numerical challenges well-known for high eigenvalues such as spectral pollution [LABV90, LS10]. Furthermore, it is known that (higher) eigenvalues may “cross” in its dependence on the random parameter, leading to difficulties in the correct identification and “matching” of corresponding eigenvalues and eigenfuctions [Kat95]. We emphasize that our method and error estimates is targeted at accurate and efficient approximations for single realizations/samples and we do not consider the possibility of crossing eigenvalues in our numerically experiments either.
5.1 Experiments for the OLOD method
We consider $\Omega=[0,1]^{d}$ for $d=1,2$. The spectral bounds are taken to be $\alpha=0.1$ and $\beta=1$. We consider $\mathfrak{k}=3$ for the LOD formulation and use the standard finite element method on a fine mesh as our reference solver. We measure the root mean square error (RMSE) between the reference eigenvlaue and the OLOD eigenvalues against various probabilities $p$ and the coarse mesh size $H$ over $N=200$ samples.
@@ 5.1.1 One-dimensional experiments
In the one-dimensional case, we use a fixed value of $\varepsilon=2^{-7}$ where we randomly assign the value $0.1$ to each interval of length $\varepsilon$. The fine mesh size is taken to be $h=2^{-8}$. Figure 3 shows the error convergence with respect to $H$ (left) and the defect probability $p$ (right). In Figure 3(a), we observe a second order convergence, which confirms our theoretical findings in Theorem 4.9. For finer $H$, we see that the convergence becomes slower the larger $p$ is. This is most probably due to the consistency error in the offline-online strategy, which is expected to grow with $p$ and to dominate LOD spatial discretization error for sufficiently fine meshes. Similar $H$-convergence rates can also be observed in the two-dimensional experiments (data not shown). Having confirmed the expected $H$-convergence, we completely focus on the error dependence on $p$ in the following and choose rather small values of $H$ to concentrate on the consistency error.
@@
Figure 3: The one-dimensional random checkerboard for $\varepsilon=2^{-7}$ with a sample size $200$.

### v2

@@ 4.3 OLOD method
where the constant $\widehat{C}_{\lambda}$ follows from a similar explanation for $\widetilde{C}_{\lambda}$ in (23). We use Lemma 2.3, Theorem 2.2, (12) of Theorem 2.7 and the stability of $\mathscr{C}$ in the energy norm together with the triangle inequality to estimate the last term as
$\!\!\!\!\!\!\!\!\!\!\!\!\norm{\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}-\widehat{m}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)$
$=m\left(\left(1-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}},\left(1-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)-m\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)$
|$\leq\absolutevalue{m(\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}},\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}})}+2\absolutevalue{m(\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})}$ |
@@
$\lesssim\left(\norm{\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+2\norm{\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\right)H^{2}$
@@
$\lesssim\Bigg{(}\norm{\left(\mathscr{C}-{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{\left(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+2\norm{\left(\mathscr{C}-{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}$
@@
$\hskip 56.9055pt+2\norm{\left(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}+3\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}\Bigg{)}H^{2}$
$\lesssim\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}\left(g^{2}(\mathfrak{k})+2g(\mathfrak{k})+3+\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)^{2}\mathfrak{k}^{d}+2\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}\right)H^{2}.$ (27)
@@
${\widetilde{a}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)-\lambda^{(l+j)}}={a\left((1-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)-\lambda^{(l+j)}}.$
Since $\left(1-{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}\in V^{\text{ms}}_{H,\mathfrak{k}}$, by the construction in (21), we obtain that
$\begin{split}a\left(\left(1-{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)&\leq\norm{\left(1-\mathscr{C}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{\left(\mathscr{C}-\mathscr{C}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\leq\norm{\left(1-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{\left(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{\left(\mathscr{C}-\mathscr{C}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}.\end{split}$
Noting that $\norm{\left(1-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}^{2}=a\left(\left(1-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}},\left(1-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)$, by combining with (20), we arrive at
${\widetilde{a}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)-\lambda^{(l+j)}}$ $\leq\norm{u^{(l+j)}-\hat{u}^{(l+j),\text{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda^{(l)}\norm{u^{(l+j)}-\hat{u}^{(l+j),\text{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}$
$\hskip 56.9055pt+\norm{\left(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{\left(\mathscr{C}-{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}$
$\lesssim\norm{u^{(l+j)}-\hat{u}^{(l+j),\text{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{\left(\mathscr{C}_{\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\norm{\left(\mathscr{C}-{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}.$ (28)
@@
Using (11), and substituting (4.3) and (28) in (26), we find the desired estimate as
$\begin{split}\left|\frac{\hat{\lambda}^{(l+j)}_{H,\mathfrak{k}}-\lambda^{(l+j)}}{\lambda^{(l+j)}}\right|&\lesssim\left(\max\limits_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}\norm{\hat{u}^{(l+j)}_{H}}_{\mathcal{A}}^{2}+\left(H^{2}+g(\mathfrak{k})+\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}\right)^{2}+\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)^{2}\mathfrak{k}^{d}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}\\ &\hskip 56.9055pt+g(\mathfrak{k})\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}+\left(3+g(\mathfrak{k})+\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}\right)H^{2}\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}\\ &\lesssim\left(H^{2}+g(\mathfrak{k})+\left(\max\limits_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}+\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)^{2}\mathfrak{k}^{d}\right)\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}^{2}.\end{split}$
The result follows, since $\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}^{2}$, similar to (23), can be bounded above by a constant due to $\widehat{a}(\cdot,\cdot)$-coercivity given in (13). ∎
@@ 5 Numerical experiments
In this section, we illustrate our theoretical developments with several numerical experiments. We focus on the OLOD method and its alternate variant with the $s$-values from Section 2.5. For details in the calculation of $E_{T}$ (if desired for error control) and discussions of the run-time complexity, we refer to [MV22, Section 5]. Note that aspects on the implementation of the LOD (without offline-online strategy) are explained in [EHMP19]. Our code is based on [HK19] and [Ver21], and freely available at [Kol25].
@@
In our numerical experiments, we will focus on the approximation of the lowest non-trivial eigenvalue, but emphasize that the theory of Section 4 is also valid for multiple and/or higher eigenvalues. By our focus on the lowest eigenvalue, we circumvent any possible numerical challenges well-known for high eigenvalues such as spectral pollution [LABV90, LS10]. Furthermore, it is known that (higher) eigenvalues may “cross” in its dependence on the random parameter, leading to difficulties in the correct identification and “matching” of corresponding eigenvalues and eigenfuctions [Kat95]. We emphasize that our method and error estimates are targeted at accurate and efficient approximations for single realizations or samples, and that we do not consider the possibility of eigenvalue crossings in our numerical experiments.
In our experiments discussed in Sections 5.1 and 5.2, we consider the domain $\Omega=[0,1]^{d}$ for $d=1,2$. The spectral bounds are set to $\alpha=0.1$ and $\beta=1$. We consider $\mathfrak{k}=3$ for the LOD formulation and use the standard finite element method on a fine mesh with the fixed mesh size $h=2^{-8}$ as our reference solver. We further set the multiscale parameter $\varepsilon=2^{-7}$. We measure the root mean square error (RMSE) between the reference eigenvalue and the OLOD eigenvalues against various probabilities $p$ and the coarse mesh size $H$ over $N=200$ samples. To compute eigenvalues, we use the built-in functions in the scipy.sparse package, which are based on the ARPACK solver. In all cases we use a tolerance $10^{-4}$.
@@ 5.1.1 One-dimensional experiments
In the one-dimensional case, we randomly assign the value $0.1$ to each interval of length $\varepsilon$. Figure 3 shows the error convergence with respect to $H$ (left) and the defect probability $p$ (right). In Figure 3(a), we observe a second order convergence, which confirms our theoretical findings in Theorem 4.9. For finer $H$, we see that the convergence becomes slower the larger $p$ is. This is most probably due to the consistency error in the offline-online strategy, which is expected to grow with $p$ and to dominate LOD spatial discretization error for sufficiently fine meshes. Similar $H$-convergence rates can also be observed in the two-dimensional experiments (data not shown). Having confirmed the expected $H$-convergence, we completely focus on the error dependence on $p$ in the following and choose rather small values of $H$ to concentrate on the consistency error.
@@
Figure 3: The one-dimensional random checkerboard.